40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Engineering group Lab elog  Not logged in ELOG logo
Message ID: 261     Entry time: Wed Jan 19 18:31:18 2022
Author: Stephen 
Type: Progress 
Category: RTS for COC 
Subject: Anneal of SN1009 at 300C for 10 hours 

SN1009 (80 mm witness sample) was annealed between 19 and 20 January 2022 - used the large furnace in Downs 221, atmosphere was air, used glass petri dishes to protect optic (from Gabriele and CRiMe). Start time was 18:28 Pacific, and sample was removed at 17:50 Pacific on the next day.

Controller parameters:

Ramp up rate = 100 °C per hour
Hold temp = 300 °C
Hold time = 10 hours
Ramp down rate = 100 °C per hour

Witness RTD observed a ~20 °C overshoot and ~13 °C offset from controller RTD. These parameters were consistent with the elog entries ENG_Labs/253 and ENG_Labs/254. See attached image of the temperature profile (.xlsx file is source document).

Optic is now with Liyuan in RTS for recharacterization after anneal.

 

Attachment 1: anneal_sn1009_300C_20220120.png  54 kB  Uploaded Tue Jan 25 09:51:31 2022  | Hide | Hide all
anneal_sn1009_300C_20220120.png
Attachment 2: anneal_sn1009_300C_20220120.xlsx  226 kB  Uploaded Tue Jan 25 09:53:53 2022
ELOG V3.1.3-